Determine the best level of output for a perfectly competitive firm that sells its 

product at P = $4 and faces TC = 0.04Q3

– 0.9Q2 + 10Q + 5. Will the firm produce this level of 

output? Why?


1
Expert's answer
2020-09-08T10:13:28-0400

Total Revenue(TR) =Price "\times" Q

TR=4Q


Profit =Total Revenue(TR) - Total Costs (TC)


Profit ="4Q\u2013(0.04Q^{3} \u20130.9Q^{2} +10Q +5)"

Profit ="\u20130.04Q^{3} +0.9Q^{2} \u20136Q\u20135"


"\frac{dP} {dQ}" ="\u20130.12Q^{2} +1.8Q \u20136"


"\frac{dP} {dQ} =0"


"\u20130.12Q^{2} +1.8Q\u20136"


"\u2013Q^{2} +15Q-50=0"


"Q(\u2013Q+10)-5(10\u2013Q) =0"


"(Q\u20135)(Q\u201310)=0"


"(Q=5) and (Q=10)"


For Q=5;Profit=

"\u20130.04(5)^{3} +0.9(5)^{2} \u20136(5)\u20135"

= –17.5

For Q=10;Profit ="\u20130.04(10)^{3} +0.9(10)^{2} \u20136(10)\u20135"

= –15


The firm would be operating at a loss hence the firm should not produce this level of output evidenced by the sell of the firm’s output at $4 per unit gives negative profits.